Fermi National Laboratory


Accelerator Update for February 16, 2005 - February 18, 2005

Wednesday February 16
The day shift began with Operations monitoring store 3986, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MCenter experiment taking beam.

At 10:12 AM, as scheduled, Pbar experts accessed their AP1 transfer line for maintenance work. RF experts began work on the Booster RF and the Main Injector (MI) low level RF system.

Operations reported that the Booster and MI RF work was completed by 10:33 AM.

Operations reestablished beam to MCenter at 10:39 AM.

The AP1 accesses ended and Pbar resumed stacking at 11:30 AM.

Operations put the Recycler into shot setup at 11:44 AM to stash antiprotons. Shot setup halts stacking and Switchyard (SY) beam.

Pbar resumed stacking at 12:58 PM.

At 1:18 PM, Operations reported seeing signs of movement in a Collider Detector at Fermilab (CDF) quadrupole magnet. CDF experimenters investigated and found that work was occurring on their collision hall air conditioning at the time.

MCenter resumed taking beam at 1:31 PM.

The Neutron Therapy Facility (NTF) requested beam at 2:02 PM.

At 7:06 PM, MCenter called about their lack of beam. Operators investigated and found the SY CAMAC link down due to a failed 5V power supply. An operator and an expert replaced the supply.

At 7:32 PM, a MI RF power supply tripped off due to a 13.8kV unbalance. Operators couldn't reset the trip and called in an expert. He found a bad fuse and made a temporary repair.

Pbar resumed stacking at 9:40 PM.

At 10:18 PM, a Booster device (MP01) had a regulation problem. Operators called in an expert. This problem held off downstream beam.

At 10:22 PM, Operations noticed a problem with the Argon gas used for the Meson SWICs. An expert suggested an adjustment.

At 11:10 PM, the same MI RF power supply that tripped off earlier tripped again. The expert was called back in. He reported that a wire he had used as a bypass had slipped off. He made a better repair.

Thursday February 17
The midnight shift began with Operations monitoring store 3986, the stack, and the Recycler stash, and with no beam elsewhere due to a Booster problem with MP01.

The Booster expert reseated a module in the low level RF system and Operations reported at 12:09 AM that beam had returned.

Pbar resumed stacking at 12:18 AM.

At 12:28 AM, B:MP01 stopped firing. An operator and an expert investigated and restored beam about ten minutes later. However, the expert continued to work on a notcher problem that held off MiniBooNE. Another expert came in to help and they fixed the problem.

MiniBooNE resumed taking beam at 1:48 AM.

Around 2:26 AM, a Booster RF station (BRF4) tripped off for the first of many times through the shift. At 6:20 AM, operators bypassed the station.

An RF expert had BRF4 back up and running by 7:23 AM.

At 8:17 AM, Operations reported that BRF4 had tripped off twice more. It reset okay.

Operations terminated store 3986 at 9 AM.

At 10:12 AM, Operations reported that there were shot setup problems with Pbar reverse protons. Operators contacted an expert who found and fixed a problem.

Operations put the TeV into shot setup at 10:13 AM.

NTF requested beam at 10:39 AM.

Operations established store 3987 at 12:55 PM with an initial luminosity of 66.3E30. The low luminosity was due to the loss of the first two antiproton transfers, which was caused by a phase problem. Pbar resumed stacking.

MCenter resumed taking beam at 1:04 PM.

MTest began taking beam at 3:41 PM.

Operations readied the Recycler to stash antiprotons at 7:06 PM.

Pbar resumed stacking at 8:15 PM.

MCenter resumed taking beam 8:16 PM.

Operations reported around 10:50 PM that they concluded a TeV flying wire study.

Friday February 18
The midnight shift began with Operations monitoring store 3987 and the stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

Operations reported at 7:35 AM that operators had search and secured the NuMI decay tunnel in preparation for their startup today.

The Plan for Friday and the Weekend
The plan for today is to terminate the store around 1 PM today and then put the TeV into shot setup. During the shot setup, the Run Coordinator will allow Linac experts a brief access. The weekend will consist of stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/18/2005   email Fermilab